### v2

# Faster Neighborhood Attention: Reducing the $\mathcal{O}(n^{2})$ Cost of Self Attention at the Threadblock Level

Abstract
Neighborhood attention [8] reduces the cost of self attention by restricting each token’s attention span to its nearest neighbors. This restriction, parameterized by a window size and dilation factor, draws a spectrum of possible attention patterns between linear projection and self attention. Neighborhood attention, and more generally sliding window attention patterns, have long been bounded by infrastructure, particularly in higher-rank spaces (2-D and 3-D), calling for the development of custom kernels, which have been limited in either functionality, or performance, if not both. In this work, we aim to massively improve upon existing infrastructure by providing two new methods for implementing neighborhood attention. We first show that neighborhood attention can be represented as a batched GEMM problem, similar to standard attention, and implement it for 1-D and 2-D neighborhood attention. These kernels on average provide 895% and 272% improvement in full precision latency compared to existing naive CUDA kernels for 1-D and 2-D neighborhood attention respectively. We find certain inherent inefficiencies in all unfused implementations of neighborhood attention that bound their performance and hinder their lower-precision scalability. Motivated by the progress made into fused dot-product attention kernels, we developed fused neighborhood attention; an adaptation of fused dot-product attention kernels that allow fine-grained control over attention across different spatial axes. Known for reducing the quadratic time complexity of self attention to a linear complexity, neighborhood attention can now enjoy a reduced and constant memory footprint, and record-breaking half precision latency. We observe that our fused implementation successfully circumvents some of the unavoidable inefficiencies in unfused implementations. While our unfused GEMM-based kernels only improve half precision performance compared to naive kernels by an average of 496% and 113% in 1-D and 2-D problems respectively, our fused kernels improve naive kernels by an average of 1607% and 581% in 1-D and 2-D problems respectively. These translate into up to 97% improvement in inference speed in existing models based on neighborhood attention, and additionally extend its applicability in vision (image and video) as well as other modalities. We will open source our kernels at github.com/SHI-Labs/NATTEN.
Figure 1: Overview of average improvement in speed on Ampere from our proposed implementation. Baseline is the set of naive CUDA kernels introduced in Neighborhood Attention Transformer [8]. GEMM-based NA improves 1-D problems by an average of 895% (FP32) and 496% (FP16), and 2-D problems by an average of 272% (FP32) and 113% (FP16). GEMM-based NA does not implement 3-D problems yet. Fused NA boosts performance further and improves 1-D problems by an average of 1072% (FP32) and 1607% (FP16), 2-D problems by an average of 381% (FP32) and 581% (FP16), and 3-D problems by an average of 460% (FP32) and 552% (FP16).
1 Introduction
Inarguably among the most highly utilized and influential primitives in modern deep learning, attention has long been cited for its complexity and memory footprint, especially when the query and context sets are identical (self attention). For years since its adoption in deep learning [20], the most common implementation of attention was through two batched GEMM (General Matrix-Matrix Multiplication) operations, sometimes referred to as “BMM-style” attention. This implementation stores attention weights to global memory, which can become a bottleneck in both speed and memory footprint. As the number of possible attention weights grow, the problem gets bounded by global memory bandwidth, as well as the amount of global memory required.
Over the past few years, some works proposed attention implementations in which attention weights are not stored to global memory, but instead kept in on-chip memory (shared memory or register file) [16, 5], which are then directly applied to corresponding values, writing the resulting attention outputs to global memory. These implementations, known as fused or memory-efficient attention, reduce the number of global memory accesses in addition to global memory usage. Thanks to the first open-source implementation, Flash Attention [5], these fused attention kernels have started replacing the standard BMM-style implementations in many deep learning frameworks and inference engines such as PyTorch [14].
Orthogonal to these efforts, many have sought to address the quadratic complexity in self attention, which is the more predominantly used operation in vision models. Neighborhood attention [8] is one such method in which each query token is restricted to only interact with its nearest neighboring context tokens. In most cases, this pattern creates a sliding window pattern, which in deep learning can also be seen in the convolution operator. This restriction can similarly be parameterized by a window size and dilation factor, and reduces the quadratic complexity of self attention down to a linear complexity. This reduction is however very difficult to implement at the tensor library or deep learning framework level. Tensor views can represent sliding window attention [17], but not the neighborhood attention pattern. In addition, GEMM operators do not typically support taking in the sliding window views in higher-rank spaces (2-D and 3-D) without explicit copying, which in practice undoes the theoretical efficiency gain from the reduced attention complexity. As a result, neighborhood attention was proposed along with an extension carrying naive CUDA kernels [8] implementing the two BMM operators. While those kernels can provide competitive FP32 performance, and in some cases even FP16/BF16 performance, they fall short of being generally adoptable in larger scale experiments. In addition, fused attention kernels such as Flash Attention effectively eliminate the $\mathcal{O}(n^{2})$ memory footprint in self attention, while also reducing latency significantly [5], making subquadratic attention patterns that are only possible to implement in “BMM-style” less feasible in practice.
In this work, we present two new classes of neighborhood attention kernels: GEMM-based BMM-style kernels (GEMM NA), and fused kernels (Fused NA), which are aimed at providing significantly improved infrastructure for neighborhood attention. We first show that neighborhood attention, and by extension sliding window attention, both of which are GEMV (General Matrix-Vector Multiplication) problems, can be expressed as GEMM problems with space-aware tiling and gather/scatter fusion. This would allow implementing such attention patterns with GEMMs, which in turn can benefit greatly from the underlying efficiency of GEMM kernels. We then extend the same logic to fused attention kernels by factoring out all logic that corresponds to the “token space”, and writing specializations that assume higher rank spaces, such as 2-D and 3-D. This, in theory, allows any fused attention kernel to be modified to accommodate token spaces of any rank. In addition, most of the logic is evaluated at compile time, resulting in minimal overhead stemming from overly complicated kernels, branch divergence, or any of the like. Finally, the structural simplicity of the resulting fused neighborhood attention kernels allows for easily adding features such as varying window size / dilation values across spatial axes, causal masking, and more.
Figure 2: Illustration of the spectrum of possible attention patterns provided by neighborhood attention. Neighborhood attention attempts to center the query token (red) within the context window (blue), as opposed to standard sliding window attention, which forces it. Neighborhood attention with window size 1 is equivalent to a linear projection (no attention). Neighborhood attention approaches self attention as window size grows, and matches it when equal to input size. Dilated windows can introduce more global context without any additional compute. Causal neighborhood attention shifts the neighborhood to end with the query.
2 Related works
Attention being adopted as a primitive is largely owed to the Transformer architecture [20], which despite its original application in machine translation rose to a predominant deep learning architecture, in addition to its design and use of the attention operator extending across applications and modalities [13, 6, 1, 15]. Attention is defined as an operation between a query set and a context set, both of which undergo linear projections, with the latter projected into a pair of key and value sets. Scaled dot product of queries and keys, $A$, is mapped into a probability distribution through the softmax operator, which produces the final attention weights, $P$. The output is a set of tokens, each derived from the weighted sum of all value tokens according to the token’s attention weights. It can be formally expressed as:
$Attention(Q,K,V)=\overbrace{softmax\left(\underbrace{\frac{QK^{T}}{\sqrt{d}}}_{A}\right)}^{P}V,$ (1)
where $\sqrt{d}$ is the scale term, and $d$ is the number of dimensions for which the dot product is computed (the embedding dimension of $Q$ and $K$). Dot product self attention, or simply, self attention, is a special case in which the query and context sets are identical. This means for a set of $n$ tokens, the attention weights matrix, $P$, is $\in\mathbb{R}^{n\times n}$, incurring an $\mathcal{O}(n^{2})$ time and space complexity. In addition, the softmax term requires reducing entire rows of the attention weight matrix, making kernel fusion much more challenging. However, fused attention kernels would at the very least eliminate the $\mathcal{O}(n^{2})$ global memory footprint, itself a significant achievement that would pave the way for the application of attention across modalities. To our knowledge, the first open-source implementation of a fused multi-headed attention (FMHA) kernel was contributed to the NVIDIA Apex  [^1] https://github.com/NVIDIA/apex project by Young-Jun Ko, which was primarily used for accelerating certain Transformer-based language models, but heavily limited in terms of supported problem sizes. Milakov and Gimelshein [12] presented a technique for computing partial softmax statistics, which when aggregation over all subsets of the original set results in exact softmax computation. This method makes the fusion of attention kernels more practical, because they would no longer be required to compute a full row of attention weights [16]. Dao et al. [5] presented and open-sourced Flash Attention, a generic fused attention kernel along with an efficient backward-pass kernel, both of which outperformed standard BMM-style attention implementations available. It was quickly adopted by many standard libraries, such as PyTorch [14], and further improved in the more recent Flash Attention V2 [4].
Parallel to these efforts, many proposed restricted attention patterns, in which the context for every query token is restricted in order to generate fewer attention weights, which in turn reduces the $\mathcal{O}(n^{2})$ time and space complexity. Stand-alone self attention (SASA) [17] is a simple 2-dimensional sliding window attention pattern, which was shown to effectively replace convolution operations in ResNet [9] variants. Noting challenges in implementing such patterns without incurring additional overhead from tensor copies and expansion, the authors later moved away from explicit sliding window attention patterns to alternatives that relaxed the sliding window movement in HaloNet [19]. In addition to these works, sliding window attention patterns in 1-dimensional spaces has been explored in language, in works such as Sparse Transformers [3], Longformer [2], BigBird [22], and more recently, Mistral [10]. Neighborhood attention [8, 7] is the practice of restricting the context of each token to its nearest neighbors, which in many cases behaves like a sliding window pattern, with the exception of corner cases in which the query cannot be centered in a sliding window. Per the original definitions from SASA [17], and Longformer [2], the sliding window can go out of bounds, in which case the attention weights corresponding to out-of-bounds context tokens are masked. This means tokens close to spatial bounds interact with fewer context tokens. This difference allows neighborhood attention to approach self attention as window size grows. In addition, neighborhood attention defines a dilation factor [7], where the number of such corner cases only increase. Fig. 2 depicts possible attention pattern for a single token under different neighborhood attention parameters. Facing similar implementation challenges as previous works, particularly SASA [17], neighborhood attention was implemented with naive CUDA kernels, and packaged as a PyTorch extension, $\mathcal{N}ATTEN$. While those kernels have accelerated research in this direction, they simply were not intended to fully utilize the underlying hardware. The only exception is the tiled kernels, which are somewhat better optimized, but only implement the $QK^{T}$ operation for 2-dimensional problems with $32$ dimensions per attention head, and window sizes up to $13\times 13$. In addition, with the rise of fused attention kernels such as Flash Attention [5], such implementations would have a hard time competing in terms of latency, and in most cases unable to compete in terms of global memory footprint.
To address these challenges, we present two new implementations and integrate them into $\mathcal{N}ATTEN$, aiming to accelerate all neighborhood attention applications, reduce their existing memory overhead, and extend existing functionality. We first simplify the operations that implement neighborhood attention’s forward and backward pass into 3 primary operators, and show each can be implemented with a batched GEMM kernel with either a fused gather or scatter operation. We then point out key limitations in unfused neighborhood attention implementations that would prevent them from achieving competitive performance compared to standard BMM-style attention implementations. Motivated by this, and the progress made in fused attention kernels, we propose fused neighborhood attention, borrowing from some of the data loading and tiling logic in our batched GEMM methodology. We implement both approaches using CUTLASS [18], and show that the batched GEMM kernels can successfully outperform most existing $\mathcal{N}ATTEN$ kernels in performance, and that our fused kernels can outperform our batched GEMM kernels while reducing the memory footprint.
3 Methodology
Figure 3: A simplified illustration of our of GEMM-based implementation of the 2-D $QK^{T}$ operation. Queries are tiled according to their original spatial rank, which in this case is 2. Keys are first restricted according to the context window size, which is just a haloed query tile. Depending on the GEMM configuration, the haloed key tile may or may not be further tiled into smaller tiles compatible with the underlying GEMM routine. Once the GEMM computes the dot product between the query and key tiles, the epilogue will scatter valid attention weights into the attention weights tensor, and discards the rest.
Herein we describe the 3 primary operations (excluding softmax) that are required to implement a full neighborhood attention forward and backward pass. We then show that each operation can be expressed as a batched GEMM problem, as long as tiling is done according to the underlying spatial rank, and attention weights are scattered and gathered. The latter is a major bottleneck for all BMM-style implementations of neighborhood attention, which in particular limits their scalability on modern hardware (i.e. Ampere and Hopper architectures.) We then introduce our fused neighborhood attention formulation which builds upon our batched GEMM formulation and tiles according to the underlying spatial rank. This approach no longer requires gathering or scattering attention weights to global memory by definition, and thereby circumvents the aforementioned bottleneck and boosts neighborhood attention performance on modern architectures.
3.1 Operators
A standard BMM-style attention forward pass (excluding softmax) is comprised of two operations: $QK^{T}$, which produces pre-softmax attention weights ($A$), and $PV$, which applies post-softmax attention weights ($P$) to values ($V$). These operations are different in that the first is a row-major by column-major GEMM, while the other is a row-major by row-major GEMM.
In the case of neighborhood attention, and sliding window attention in general, these will become GEMV problems. In $QK^{T}$, each query token (vector) is multiplied by its neighboring or surrounding key tokens (matrix), and in $PV$, the set of attention weights corresponding to each query token (vector) is multiplied by corresponding value tokens (matrix). We dub the $QK^{T}$ operation “Pointwise-Neighborhood” (PN) and the $PV$ operation “Neighborhood-Neighborhood” (NN), because the same operations can be reused in the backward pass. PN can compute the gradient for post-softmax attention weights ($\nabla P$) when it is fed the output gradient and $V$ instead of $Q$ and $K$. Similarly, NN can compute the gradient for $Q$ ($\nabla Q$), if it is input the pre-softmax attention gradient ($\nabla A$) and $K$ instead of $A$ and $V$. We define a third operator, which can compute gradients for both $K$ and $V$ given the correct input operands: Inverse-Neighborhood (IN). This operation is mostly similar to NN, but differs in the way attention weights are gathered, as well as the number of attention weights. IN may also require loading more attention weights for every token, because unlike in sliding window attention, the relationship between query and context tokens in neighborhood attention is not commutative. In other words, query token at coordinate $i$ attending to context token at coordinate $j$ does not imply that query token at coordinate $j$ attends to context token at coordinate $i$.
BMM-style implementations of standard self attention have a clear edge over neighborhood and sliding window attention implementations, because they are GEMM problems and by extension not as bound by memory as the latter, both of which are GEMV problems. In addition, GEMV problems cannot target specialized matrix multiply-and-accumulate (MMA) cores, such as Tensor Cores. We aim to resolve this issue by formulating all three operators as GEMM problems, in order to better utilize modern hardware accelerators.
3.2 Batched GEMM NA
Mapping GEMV problems to GEMM problems is straightforward: A subset of individual vectors (a tile at the threadblock level) can be packed into a matrix, and their corresponding individual matrices can be concatenated. Elements in the resulting matrix will have to be scattered according to the original GEMV problem. This approach is not practical unless the sets of column vectors from the individual matrices intersect, which, when eliminated, will reduce the GEMM problem size. In the context of neighborhood and sliding window attention, the aforementioned individual matrices are context tokens, all of which have coordinates in the original token space, and are selected according to the query coordinate. Due to this, as long as query vectors that are packed into a matrix are spatially proximate, their corresponding context tokens will overlap greatly, which reduces the GEMM problem size. We therefore propose modifying a standard batched GEMM kernel as follows in order to implement the above mentioned operators (PN, NN, and IN):
  1. 1.
    GEMM tile sizes corresponding to number of tokens (rows and columns) are mapped to multi-dimensional shapes according to the input spatial rank. For example, a GEMM with $64$ rows is mapped to an $8\times 8$ tile for 2-dimensional problems, or to a $4\times 4\times 4$ tile for 3-dimensional problems, but the exact tile shape can either be fixed, or chosen according to problem size.
  2. 2.
    Global pointer iterators and accessors are modified to iterate according to the original layout in global memory instead of a standard row or column major matrix layout.
  3. 3.
    Attention weights are only scattered to and gathered from global memory, and cannot be accessed as vectors. This especially impacts kernels targeting Tensor Cores, where asynchronous data prefetching becomes impossible.
Figure 4: A simplified illustration of fused neighborhood attention. Queries, and the key-value pair are tiled according to their spatial rank (1-D, 2-D, 3-D), with the latter haloed to include the entire neighborhood for all corresponding queries in the query tile. The haloed KV tile is further optionally tiled into sub-tiles according to the underlying threadblock-scope GEMM shape. Resulting attention weights from the first GEMM are masked according to neighborhood attention parameters, before undergoing online softmax scaling, and going through the second GEMM with the corresponding value sub-tile.
We implemented these concepts by modifying the implicit GEMM kernels in CUTLASS (2.X API) which implement convolution, and successfully implemented 1-dimensional and 2-dimensional kernels that compute the three neighborhood attention operators. Fig. 3 shows an illustration of the GEMM-based PN kernel. 2223-dimensional kernels will be implemented in the future. The first change is relatively inexpensive, but the second change incurs additional indexing logic that may contribute to latency. The final change is a major bottleneck, which we later show cause lower-precision kernels (FP16/BF16) to regress, to the point of providing little to no improvement in latency compared to their full precision (FP32/TF32) counterparts. NN and IN suffer from this issue more significantly, because gathering attention weights by definition does not allow asynchronously prefetching data for one operand (attention weights) in the GEMM mainloop unless the scalar data type meets the minimum access size (32 bits on Ampere). This forces our FP16/BF16 GEMM-based kernels to fall back to standard global reads, which significantly impact achievable latency. To our knowledge, this issue is unavoidable in most cases, and will continue to be a bottleneck as long as attention weights are stored in global memory. Every attention weight corresponds to a query and context coordinate, and even in theory, vectorized access cannot replace the gather/scatter of attention weights unless: 1. window size is known at compile time (which is not very practical especially in 2-D and 3-D problems), and 2. window size is a power of 2 (which it never is in neighborhood attention because window size must be an odd number).
3.3 Fused NA
We extend our methodology for implementing neighborhood attention operators using batched GEMM kernels to fused attention kernels like Flash Attention [5]. This is motivated by not only the potential to reduce latency and memory footprint, but also to circumvent the bottleneck in the batched GEMM and naive kernels, scattering and gathering attention weights to global memory. Since attention weights are only computed at the threadblock level and never stored in global memory in fused kernels, the bottleneck will simply cease to exist. We started off with FMHA (developed by the xFormers [^3] https://github.com/facebookresearch/xformers/ team), a fused multi-headed attention kernel based on the CUTLASS 2.X API, which can target architectures even older than Ampere (Maxwell, SM50; Volta, SM70; and Turing, SM75.) By carefully applying our methodology for space-aware tiling, neighborhood attention masking, and customized pointer iterators and accessors, we successfully implemented neighborhood attention for 1-D, 2-D, and 3-D problems.  [^4] Only forward pass is implemented in this version. Fig. 4 presents an overview of how our fused kernels function when dealing with multi-axis data.
3.4 Dilation and causal masking
We implement dilation support through partitioned tensor views. A dilated neighborhood attention problem can be mapped to a set of non-dilated neighborhood attention problems over non-overlapping partitions of the input. All sub-problems can be compute within the same kernel call. We additionally define and implement causal neighborhood attention into our fused kernel, which can be crucial to certain applications where only one spatial dimension requires causal masking (i.e. video embeddings may benefit from causally masked attention across the time axis and standard attention across height and width axes.)
3.5 Auto-tuning
GEMM kernels are, among other settings, parameterized by their tiling shape. Multi-dimensional variants (2-D and 3-D neighborhood attention) can also be parameterized by their fine-grained tile sizes, introduced by our formulation. As mentioned earlier, a GEMM with row tile size 64 can be mapped to a number of valid 2-D and 3-D tiles (i.e. $x\times y$ for all positive integers $x$ and $y$ where $xy=64$, and $x\times y\times z$ for all positive integers $x$, $y$, and $z$ where $xyz=64$.) As a result, selecting tiling sizes based on factors such as problem size, hardware, and environment can further decrease achievable latency. We therefore implement a very simple auto-tuning method as a proof of concept, which is inspired by similar techniques used in packages such as cuDNN. When auto-tuning is activated, a hash map is created and maintained for the lifetime of the application, which maps problems (defined by problem size, data type, and other such factors) to a tiling configuration. Upon encountering problems for the first time, they are benchmarked with a pre-defined set of tiling configurations, which determines and the best configuration, and stores it in the hash map.
3.6 Limitations
Our formulation of GEMM-based and fused neighborhood attention kernels poses a critical question: how much overhead can one expect from switching from a standard self attention kernel to neighborhood attention? As pointed out in Sec. 3.2, our GEMM-based kernels suffer from a major bottleneck, especially in lower-precision, which stems from the scatter / gather of attention weights, which limits the ability to hide MMA prefetching latency. We consider this to be an unavoidable issue in GEMM-based neighborhood and sliding window attention kernels. Unsurprisingly, our proposed changes to fused implementations are also not free. Changes that we find unavoidable, which in some cases can cause our fused kernels to incur higher latency than the self-attention baseline (FMHA) are the following (ordered by most significant to least significant):
  1. 1.
    Kernels specialized for 2-D and 3-D problems fall back to less optimal data iterators which cannot rely on pre-computed pointer offsets as in 1-D problems, and will therefore need to both compute and apply pointer offsets per access. While this is similar to standard practice for data permutation / scattering on the fly in CUTLASS 2.X GEMMs, it is typically less performant than pre-computed offsets. In addition, computing pointer offsets incurs more integer ops with more dimensions, which is repeated per global memory access. We find this to be the most significant contributor to additional latency in our fused kernels, when compared to the baseline fused self attention kernel, FMHA. However, this is usually only significant when window size is approximately equal to input size. Smaller window sizes hide the additional latency by computing fewer attention weights, which reduces both the number of global memory accesses and MMAs.
  2. 2.
    The attention masking logic, which depends on corresponding query and context token coordinates, spatial size, and window size, introduces additional integer ops in order to map linear indices to coordinates (unlike in 1-D problems where the mapping is the identity function), which again gets more complicated with more dimensions. This, along with additional statements in the masking condition, contributes to latency, and is expected to worsen with more dimensions. We suspect that it is possible to predicate some of the attention masking logic which would reduce some of the latency, but we leave that for future efforts.
Despite these issues, we find that our fused kernels can still match or outperform our self attention baseline in approximately 97.8% of 1-D, 85.5% of 2-D, and 92.2% of all problem sizes that we benchmarked.
4 Experiments
NA Kernel % of problems matched or outperformed
neighborhood attn self attn
Naive GEMM Fused FMHA FAv2
1-dimensional neighborhood attention
Naive - 4.7 % 0.0 % 20.3 % 7.0 %
GEMM 96.4 % - 0.0 % 61.9 % 33.4 %
Fused 100.0 % 100.0 % - 97.8 % 89.8 %
2-dimensional neighborhood attention
Naive - 34.3 % 8.4 % 38.3 % 17.5 %
GEMM 65.9 % - 0.0 % 54.1 % 23.3 %
Fused 93.7 % 100.0 % - 85.5 % 59.7 %
3-dimensional neighborhood attention
Naive - - 0.3 % 49.3 % 24.3 %
Fused 99.8 % - - 92.3 % 69.4 %
Table 1: Overview of our FP16 performance benchmark. We benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in half precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers’ FMHA (our baseline) and Flash Attention V2. Note that we have not yet implemented a GEMM kernel for 3-dimensional neighborhood attention.
NA Kernel % of problems matched or outperformed
neighborhood attn self attn
Naive GEMM Fused FMHA
1-dimensional neighborhood attention
Naive - 0.1 % 0.0 % 31.2 %
GEMM 99.8 % - 29.9 % 89.4 %
Fused 100.0 % 72.1 % - 94.7 %
2-dimensional neighborhood attention
Naive - 24.6 % 18.9 % 58.1 %
GEMM 76.0 % - 36.1 % 85.5 %
Fused 81.9 % 66.8 % - 92.9 %
3-dimensional neighborhood attention
Naive - - 0.2 % 62.9 %
Fused 99.8 % - - 95.2 %
Table 2: Overview of our FP32 performance benchmark. As in Tab. 1, we benchmark naive neighborhood attention kernels against our proposed GEMM and fused kernels in full precision, over a large set of problem sizes varying in batch size, spatial size, number of attention heads, and dimensions per head, and over different window sizes and dilation values. For every problem size, we also benchmarked self attention running with the xFormers’ FMHA (our baseline). Flash Attention V2 is excluded as it does not support full precision. Note that we have not yet implemented a GEMM kernel for 3-dimensional neighborhood attention.
We evaluate the performance of our proposed methods by measuring their latency against existing kernels in $\mathcal{N}ATTEN$. Most use cases in $\mathcal{N}ATTEN$ target naive CUDA kernels, with the exception of 2-D neighborhood attention with 32-dimensional attention heads. $\mathcal{N}ATTEN$ implements tiled kernels for those cases for up to and including window size 13 × 13, and only for the $QK$ operation. However, we treat all kernels in $\mathcal{N}ATTEN$ as our baseline, and will refer to them as naive kernels hereinafter. We run a fixed set of problem sizes that vary in batch size, spatial size, number of attention heads, and dimensions per attention head through every implementation on an NVIDIA A100 GPU and measure their latency. We iterate through multiple neighborhood attention window sizes and dilation values for every problem size. A summary of these benchmarks is presented in Tab. 1 (FP16) and Tab. 2 (FP32). We find that our GEMM-based kernels can improve or match the naive latency in approximately 96% of 1-D problems (of 6150), and 66% of 2-D problems (of 5676) in half precision, and approximately 100% of the 1-D problems and 76% of the 2-D problems in full precision. Note that over 40% of the 2-D problems target tiled kernels in $\mathcal{N}ATTEN$, which we find can sometimes outperform our GEMM-based kernels. Upon re-evaluating our benchmark without the tiled kernels in $\mathcal{N}ATTEN$, our GEMM-based kernels can improve or match 84% of the 2-D problems in half precision and 100% of them in full precision. Another point of disadvantage in our GEMM-based kernels is the memory alignment issue noted in Sec. 3.2, which is especially worsened by both lower-precision and higher spatial rank (i.e. 2-D over 1-D.) On the other hand, our fused kernels improve or match the naive latency in approximately 100% of both 1-D (of 6150) and 3-D problems (of 2448) in both half precision and full precision, while only improving approximately 94% of 2-D problems in half precision and 82% of them in full precision. We again find that if we only compare to non-tiled kernels, fused neighborhood attention kernels improve or match 100% of 2-D problems in both half and full precision. We also find that our fused kernels match or outperform our GEMM kernels in 100% of both 1-D and 2-D problems in half precision, while only doing so in approximately 72% of 1-D problems and 67% of 2-D problems in full precision. Our primary takeaway from this is that despite the mentioned bottleneck in our GEMM kernels, they can still be a useful choice for problems that are heavily compute-bound, which is sometimes the case in full precision. In both Tab. 1 and Tab. 2 we also inspect the percentage of problem sizes in which using our fused neighborhood attention kernel is outperformed by the FMHA kernel. This is only to inspect additional overhead caused by our implementation, which we expect to be more noticeable in 2-D and 3-D problems. Some of the overhead may be avoidable, but our takeaway is that it is unlikely to be fully avoidable, as pointed out in Sec. 3.6.
We further present a breakdown of our benchmarks in Tab. 3 and Tab. 4, where we group cases by window size, which is the primary contributor to difference in FLOPs when considering similar input sizes. We report the average, minimum, and maximum improvement observed from switching from naive to GEMM-based, naive to fused, and GEMM-based to fused kernels. GEMM-based kernels exhibit strong performance compared to both naive and fused kernels in full precision, where fused kernels only have a very minimal edge on average. GEMM-based kernels also outperform naive kernels in half precision, especially in cases where tiled kernels are not available. While tiled kernels are sometimes the better choice, we note that they simply cannot generalize to all problem sizes as our GEMM-based kernels can, nor are they easily extensible. In addition, our auto-tuner can be extended to include BMM-style implementations as well as fused kernels, and pick the best between GEMM-based and naive (naive or tiled) according to problem size, hardware, and other such performance determining factors.
Window GEMM over naive Fused over naive Fused over GEMM
size Average Min Max Average Min Max Average Min Max
1-dimensional neighborhood attention
9 $\boldsymbol{\uparrow}$ 199 % $\boldsymbol{\downarrow}$59 % $\boldsymbol{\uparrow}$ 673 % $\boldsymbol{\uparrow}$ 705 % $\boldsymbol{\uparrow}$ 17 % $\boldsymbol{\uparrow}$ 1660 % $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\uparrow}$ 44 % $\boldsymbol{\uparrow}$ 300 %
15 $\boldsymbol{\uparrow}$ 258 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 929 % $\boldsymbol{\uparrow}$ 844 % $\boldsymbol{\uparrow}$ 57 % $\boldsymbol{\uparrow}$ 2000 % $\boldsymbol{\uparrow}$ 181 % $\boldsymbol{\uparrow}$ 49 % $\boldsymbol{\uparrow}$ 280 %
31 $\boldsymbol{\uparrow}$ 412 % $\boldsymbol{\downarrow}$12 % $\boldsymbol{\uparrow}$ 1567 % $\boldsymbol{\uparrow}$ 1217 % $\boldsymbol{\uparrow}$ 150 % $\boldsymbol{\uparrow}$ 3206 % $\boldsymbol{\uparrow}$ 183 % $\boldsymbol{\uparrow}$ 62 % $\boldsymbol{\uparrow}$ 280 %
45 $\boldsymbol{\uparrow}$ 356 % $\boldsymbol{\downarrow}$22 % $\boldsymbol{\uparrow}$ 1429 % $\boldsymbol{\uparrow}$ 1100 % $\boldsymbol{\uparrow}$ 133 % $\boldsymbol{\uparrow}$ 3150 % $\boldsymbol{\uparrow}$ 190 % $\boldsymbol{\uparrow}$ 77 % $\boldsymbol{\uparrow}$ 320 %
63 $\boldsymbol{\uparrow}$ 427 % $\boldsymbol{\downarrow}$28 % $\boldsymbol{\uparrow}$ 1697 % $\boldsymbol{\uparrow}$ 1275 % $\boldsymbol{\uparrow}$ 117 % $\boldsymbol{\uparrow}$ 3981 % $\boldsymbol{\uparrow}$ 191 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 320 %
127 $\boldsymbol{\uparrow}$ 545 % $\boldsymbol{\downarrow}$21 % $\boldsymbol{\uparrow}$ 2430 % $\boldsymbol{\uparrow}$ 1497 % $\boldsymbol{\uparrow}$ 73 % $\boldsymbol{\uparrow}$ 7494 % $\boldsymbol{\uparrow}$ 167 % $\boldsymbol{\uparrow}$ 62 % $\boldsymbol{\uparrow}$ 360 %
255 $\boldsymbol{\uparrow}$ 779 % $\boldsymbol{\downarrow}$11 % $\boldsymbol{\uparrow}$ 2598 % $\boldsymbol{\uparrow}$ 2231 % $\boldsymbol{\uparrow}$ 164 % $\boldsymbol{\uparrow}$ 8797 % $\boldsymbol{\uparrow}$ 171 % $\boldsymbol{\uparrow}$ 94 % $\boldsymbol{\uparrow}$ 291 %
511 $\boldsymbol{\uparrow}$ 1029 % $\boldsymbol{\uparrow}$ 23 % $\boldsymbol{\uparrow}$ 2781 % $\boldsymbol{\uparrow}$ 3369 % $\boldsymbol{\uparrow}$ 247 % $\boldsymbol{\uparrow}$ 10649 % $\boldsymbol{\uparrow}$ 207 % $\boldsymbol{\uparrow}$ 104 % $\boldsymbol{\uparrow}$ 344 %
767 $\boldsymbol{\uparrow}$ 1193 % $\boldsymbol{\uparrow}$ 59 % $\boldsymbol{\uparrow}$ 2473 % $\boldsymbol{\uparrow}$ 4278 % $\boldsymbol{\uparrow}$ 243 % $\boldsymbol{\uparrow}$ 10226 % $\boldsymbol{\uparrow}$ 241 % $\boldsymbol{\uparrow}$ 111 % $\boldsymbol{\uparrow}$ 386 %
1023 $\boldsymbol{\uparrow}$ 1377 % $\boldsymbol{\uparrow}$ 101 % $\boldsymbol{\uparrow}$ 2783 % $\boldsymbol{\uparrow}$ 5030 % $\boldsymbol{\uparrow}$ 444 % $\boldsymbol{\uparrow}$ 11082 % $\boldsymbol{\uparrow}$ 255 % $\boldsymbol{\uparrow}$ 155 % $\boldsymbol{\uparrow}$ 359 %
2047 $\boldsymbol{\uparrow}$ 1342 % $\boldsymbol{\uparrow}$ 181 % $\boldsymbol{\uparrow}$ 2665 % $\boldsymbol{\uparrow}$ 5721 % $\boldsymbol{\uparrow}$ 765 % $\boldsymbol{\uparrow}$ 11540 % $\boldsymbol{\uparrow}$ 322 % $\boldsymbol{\uparrow}$ 159 % $\boldsymbol{\uparrow}$ 444 %
4095 $\boldsymbol{\uparrow}$ 1434 % $\boldsymbol{\uparrow}$ 203 % $\boldsymbol{\uparrow}$ 2787 % $\boldsymbol{\uparrow}$ 6027 % $\boldsymbol{\uparrow}$ 1223 % $\boldsymbol{\uparrow}$ 11093 % $\boldsymbol{\uparrow}$ 323 % $\boldsymbol{\uparrow}$ 210 % $\boldsymbol{\uparrow}$ 453 %
All $\boldsymbol{\uparrow}$ 496 % $\boldsymbol{\downarrow}$59 % $\boldsymbol{\uparrow}$ 2787 % $\boldsymbol{\uparrow}$ 1607 % $\boldsymbol{\uparrow}$ 17 % $\boldsymbol{\uparrow}$ 11540 % $\boldsymbol{\uparrow}$ 190 % $\boldsymbol{\uparrow}$ 44 % $\boldsymbol{\uparrow}$ 453 %
2-dimensional neighborhood attention
3 × 3 $\boldsymbol{\uparrow}$ 27 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 258 % $\boldsymbol{\uparrow}$ 203 % $\boldsymbol{\downarrow}$43 % $\boldsymbol{\uparrow}$ 900 % $\boldsymbol{\uparrow}$ 164 % $\boldsymbol{\uparrow}$ 18 % $\boldsymbol{\uparrow}$ 764 %
5 × 5 $\boldsymbol{\uparrow}$ 101 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 554 % $\boldsymbol{\uparrow}$ 420 % $\boldsymbol{\downarrow}$25 % $\boldsymbol{\uparrow}$ 1990 % $\boldsymbol{\uparrow}$ 180 % $\boldsymbol{\uparrow}$ 24 % $\boldsymbol{\uparrow}$ 862 %
7 × 7 $\boldsymbol{\uparrow}$ 108 % $\boldsymbol{\downarrow}$61 % $\boldsymbol{\uparrow}$ 570 % $\boldsymbol{\uparrow}$ 509 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 3086 % $\boldsymbol{\uparrow}$ 211 % $\boldsymbol{\uparrow}$ 15 % $\boldsymbol{\uparrow}$ 1032 %
9 × 9 $\boldsymbol{\uparrow}$ 129 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 444 % $\boldsymbol{\uparrow}$ 505 % $\boldsymbol{\downarrow}$20 % $\boldsymbol{\uparrow}$ 1772 % $\boldsymbol{\uparrow}$ 163 % $\boldsymbol{\uparrow}$ 38 % $\boldsymbol{\uparrow}$ 434 %
11 × 11 $\boldsymbol{\uparrow}$ 132 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 531 % $\boldsymbol{\uparrow}$ 520 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 2298 % $\boldsymbol{\uparrow}$ 167 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\uparrow}$ 437 %
13 × 13 $\boldsymbol{\uparrow}$ 121 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 504 % $\boldsymbol{\uparrow}$ 697 % $\boldsymbol{\downarrow}$30 % $\boldsymbol{\uparrow}$ 3037 % $\boldsymbol{\uparrow}$ 252 % $\boldsymbol{\uparrow}$ 37 % $\boldsymbol{\uparrow}$ 593 %
15 × 15 $\boldsymbol{\uparrow}$ 207 % $\boldsymbol{\downarrow}$8 % $\boldsymbol{\uparrow}$ 740 % $\boldsymbol{\uparrow}$ 1082 % $\boldsymbol{\uparrow}$ 79 % $\boldsymbol{\uparrow}$ 5086 % $\boldsymbol{\uparrow}$ 269 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 618 %
17 × 17 $\boldsymbol{\uparrow}$ 251 % $\boldsymbol{\uparrow}$ 20 % $\boldsymbol{\uparrow}$ 732 % $\boldsymbol{\uparrow}$ 883 % $\boldsymbol{\uparrow}$ 106 % $\boldsymbol{\uparrow}$ 2384 % $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\uparrow}$ 53 % $\boldsymbol{\uparrow}$ 284 %
33 × 33 $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\downarrow}$13 % $\boldsymbol{\uparrow}$ 372 % $\boldsymbol{\uparrow}$ 1686 % $\boldsymbol{\uparrow}$ 469 % $\boldsymbol{\uparrow}$ 3006 % $\boldsymbol{\uparrow}$ 558 % $\boldsymbol{\uparrow}$ 429 % $\boldsymbol{\uparrow}$ 675 %
41 × 41 $\boldsymbol{\uparrow}$ 208 % $\boldsymbol{\downarrow}$8 % $\boldsymbol{\uparrow}$ 426 % $\boldsymbol{\uparrow}$ 1969 % $\boldsymbol{\uparrow}$ 588 % $\boldsymbol{\uparrow}$ 3421 % $\boldsymbol{\uparrow}$ 591 % $\boldsymbol{\uparrow}$ 400 % $\boldsymbol{\uparrow}$ 724 %
49 × 49 $\boldsymbol{\uparrow}$ 210 % $\boldsymbol{\downarrow}$12 % $\boldsymbol{\uparrow}$ 427 % $\boldsymbol{\uparrow}$ 2357 % $\boldsymbol{\uparrow}$ 753 % $\boldsymbol{\uparrow}$ 4081 % $\boldsymbol{\uparrow}$ 718 % $\boldsymbol{\uparrow}$ 468 % $\boldsymbol{\uparrow}$ 874 %
63 × 63 $\boldsymbol{\uparrow}$ 246 % $\boldsymbol{\uparrow}$ 5 % $\boldsymbol{\uparrow}$ 490 % $\boldsymbol{\uparrow}$ 3165 % $\boldsymbol{\uparrow}$ 922 % $\boldsymbol{\uparrow}$ 5510 % $\boldsymbol{\uparrow}$ 872 % $\boldsymbol{\uparrow}$ 419 % $\boldsymbol{\uparrow}$ 1096 %
All $\boldsymbol{\uparrow}$ 113 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 740 % $\boldsymbol{\uparrow}$ 581 % $\boldsymbol{\downarrow}$43 % $\boldsymbol{\uparrow}$ 5510 % $\boldsymbol{\uparrow}$ 213 % $\boldsymbol{\uparrow}$ 15 % $\boldsymbol{\uparrow}$ 1096 %
3-dimensional neighborhood attention
3 × 3 × 3 - - - $\boldsymbol{\uparrow}$ 341 % 0 % $\boldsymbol{\uparrow}$ 1651 % - - -
5 × 5 × 5 - - - $\boldsymbol{\uparrow}$ 318 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 1932 % - - -
7 × 7 × 7 - - - $\boldsymbol{\uparrow}$ 667 % $\boldsymbol{\uparrow}$ 42 % $\boldsymbol{\uparrow}$ 3583 % - - -
9 × 9 × 9 - - - $\boldsymbol{\uparrow}$ 602 % $\boldsymbol{\uparrow}$ 98 % $\boldsymbol{\uparrow}$ 1550 % - - -
11 × 11 × 11 - - - $\boldsymbol{\uparrow}$ 878 % $\boldsymbol{\uparrow}$ 160 % $\boldsymbol{\uparrow}$ 2229 % - - -
13 × 13 × 13 - - - $\boldsymbol{\uparrow}$ 1208 % $\boldsymbol{\uparrow}$ 237 % $\boldsymbol{\uparrow}$ 3130 % - - -
15 × 15 × 15 - - - $\boldsymbol{\uparrow}$ 1600 % $\boldsymbol{\uparrow}$ 241 % $\boldsymbol{\uparrow}$ 4376 % - - -
17 × 17 × 17 - - - $\boldsymbol{\uparrow}$ 822 % $\boldsymbol{\uparrow}$ 303 % $\boldsymbol{\uparrow}$ 1473 % - - -
All - - - $\boldsymbol{\uparrow}$ 552 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 4376 % - - -
Table 3: Breakdown of improvement over naive kernels (FP16). Both GEMM-based and fused NA improve the baseline naive kernels on average. However, as illustrated by the minimum improvement, there are many cases in which naive kernels may be preferable to GEMM-based, which fails to effectively hide data prefetching latency. Fused NA on the other hand manages to improve upon most cases, with the exception of many cases that target the tiled kernels in NATTEN (2-dimensional, with window size up to 13 × 13.) Very few cases ($\approx$ 0.3 %) in 3-dimensional problems where fused NA can’t always outperform naive.
Window GEMM over naive Fused over naive Fused over GEMM
size Average Min Max Average Min Max Average Min Max
1-dimensional neighborhood attention
9 $\boldsymbol{\uparrow}$ 473 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 1046 % $\boldsymbol{\uparrow}$ 639 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 1682 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$48 % $\boldsymbol{\uparrow}$ 109 %
15 $\boldsymbol{\uparrow}$ 584 % $\boldsymbol{\downarrow}$33 % $\boldsymbol{\uparrow}$ 1462 % $\boldsymbol{\uparrow}$ 739 % $\boldsymbol{\uparrow}$ 11 % $\boldsymbol{\uparrow}$ 1718 % $\boldsymbol{\uparrow}$ 27 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 100 %
31 $\boldsymbol{\uparrow}$ 829 % $\boldsymbol{\uparrow}$ 106 % $\boldsymbol{\uparrow}$ 2192 % $\boldsymbol{\uparrow}$ 1002 % $\boldsymbol{\uparrow}$ 230 % $\boldsymbol{\uparrow}$ 2247 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\downarrow}$38 % $\boldsymbol{\uparrow}$ 118 %
45 $\boldsymbol{\uparrow}$ 678 % $\boldsymbol{\uparrow}$ 100 % $\boldsymbol{\uparrow}$ 1925 % $\boldsymbol{\uparrow}$ 803 % $\boldsymbol{\uparrow}$ 182 % $\boldsymbol{\uparrow}$ 1981 % $\boldsymbol{\uparrow}$ 22 % $\boldsymbol{\downarrow}$37 % $\boldsymbol{\uparrow}$ 108 %
63 $\boldsymbol{\uparrow}$ 749 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 2118 % $\boldsymbol{\uparrow}$ 906 % $\boldsymbol{\uparrow}$ 109 % $\boldsymbol{\uparrow}$ 2567 % $\boldsymbol{\uparrow}$ 24 % $\boldsymbol{\downarrow}$35 % $\boldsymbol{\uparrow}$ 117 %
127 $\boldsymbol{\uparrow}$ 874 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\uparrow}$ 3139 % $\boldsymbol{\uparrow}$ 934 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\uparrow}$ 4179 % $\boldsymbol{\uparrow}$ 6 % $\boldsymbol{\downarrow}$39 % $\boldsymbol{\uparrow}$ 95 %
255 $\boldsymbol{\uparrow}$ 1190 % $\boldsymbol{\uparrow}$ 52 % $\boldsymbol{\uparrow}$ 3237 % $\boldsymbol{\uparrow}$ 1286 % $\boldsymbol{\uparrow}$ 45 % $\boldsymbol{\uparrow}$ 4767 % 0 % $\boldsymbol{\downarrow}$38 % $\boldsymbol{\uparrow}$ 69 %
511 $\boldsymbol{\uparrow}$ 1681 % $\boldsymbol{\uparrow}$ 196 % $\boldsymbol{\uparrow}$ 3404 % $\boldsymbol{\uparrow}$ 1894 % $\boldsymbol{\uparrow}$ 86 % $\boldsymbol{\uparrow}$ 4794 % $\boldsymbol{\uparrow}$ 5 % $\boldsymbol{\downarrow}$50 % $\boldsymbol{\uparrow}$ 67 %
767 $\boldsymbol{\uparrow}$ 2016 % $\boldsymbol{\uparrow}$ 475 % $\boldsymbol{\uparrow}$ 3276 % $\boldsymbol{\uparrow}$ 2429 % $\boldsymbol{\uparrow}$ 229 % $\boldsymbol{\uparrow}$ 4623 % $\boldsymbol{\uparrow}$ 16 % $\boldsymbol{\downarrow}$51 % $\boldsymbol{\uparrow}$ 75 %
1023 $\boldsymbol{\uparrow}$ 2209 % $\boldsymbol{\uparrow}$ 649 % $\boldsymbol{\uparrow}$ 3488 % $\boldsymbol{\uparrow}$ 2638 % $\boldsymbol{\uparrow}$ 288 % $\boldsymbol{\uparrow}$ 4668 % $\boldsymbol{\uparrow}$ 16 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 65 %
2047 $\boldsymbol{\uparrow}$ 2265 % $\boldsymbol{\uparrow}$ 888 % $\boldsymbol{\uparrow}$ 3458 % $\boldsymbol{\uparrow}$ 2923 % $\boldsymbol{\uparrow}$ 618 % $\boldsymbol{\uparrow}$ 4656 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\downarrow}$34 % $\boldsymbol{\uparrow}$ 74 %
4095 $\boldsymbol{\uparrow}$ 2319 % $\boldsymbol{\uparrow}$ 1125 % $\boldsymbol{\uparrow}$ 3546 % $\boldsymbol{\uparrow}$ 3009 % $\boldsymbol{\uparrow}$ 1056 % $\boldsymbol{\uparrow}$ 4679 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$9 % $\boldsymbol{\uparrow}$ 75 %
All $\boldsymbol{\uparrow}$ 895 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 3546 % $\boldsymbol{\uparrow}$ 1072 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 4794 % $\boldsymbol{\uparrow}$ 21 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 118 %
2-dimensional neighborhood attention
3 × 3 $\boldsymbol{\uparrow}$ 135 % $\boldsymbol{\downarrow}$58 % $\boldsymbol{\uparrow}$ 607 % $\boldsymbol{\uparrow}$ 161 % $\boldsymbol{\downarrow}$65 % $\boldsymbol{\uparrow}$ 810 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 301 %
5 × 5 $\boldsymbol{\uparrow}$ 260 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 1021 % $\boldsymbol{\uparrow}$ 290 % $\boldsymbol{\downarrow}$62 % $\boldsymbol{\uparrow}$ 1301 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\downarrow}$51 % $\boldsymbol{\uparrow}$ 348 %
7 × 7 $\boldsymbol{\uparrow}$ 246 % $\boldsymbol{\downarrow}$47 % $\boldsymbol{\uparrow}$ 1006 % $\boldsymbol{\uparrow}$ 331 % $\boldsymbol{\downarrow}$67 % $\boldsymbol{\uparrow}$ 2005 % $\boldsymbol{\uparrow}$ 45 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 436 %
9 × 9 $\boldsymbol{\uparrow}$ 238 % $\boldsymbol{\downarrow}$32 % $\boldsymbol{\uparrow}$ 648 % $\boldsymbol{\uparrow}$ 303 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 1052 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\downarrow}$46 % $\boldsymbol{\uparrow}$ 156 %
11 × 11 $\boldsymbol{\uparrow}$ 244 % $\boldsymbol{\downarrow}$29 % $\boldsymbol{\uparrow}$ 856 % $\boldsymbol{\uparrow}$ 294 % $\boldsymbol{\downarrow}$61 % $\boldsymbol{\uparrow}$ 1459 % $\boldsymbol{\uparrow}$ 20 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 144 %
13 × 13 $\boldsymbol{\uparrow}$ 220 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 789 % $\boldsymbol{\uparrow}$ 393 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 1835 % $\boldsymbol{\uparrow}$ 59 % $\boldsymbol{\downarrow}$50 % $\boldsymbol{\uparrow}$ 217 %
15 × 15 $\boldsymbol{\uparrow}$ 401 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 1175 % $\boldsymbol{\uparrow}$ 698 % 0 % $\boldsymbol{\uparrow}$ 2845 % $\boldsymbol{\uparrow}$ 56 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 240 %
17 × 17 $\boldsymbol{\uparrow}$ 510 % $\boldsymbol{\uparrow}$ 90 % $\boldsymbol{\uparrow}$ 1246 % $\boldsymbol{\uparrow}$ 565 % $\boldsymbol{\uparrow}$ 23 % $\boldsymbol{\uparrow}$ 1347 % $\boldsymbol{\uparrow}$ 10 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 73 %
33 × 33 $\boldsymbol{\uparrow}$ 860 % $\boldsymbol{\uparrow}$ 308 % $\boldsymbol{\uparrow}$ 1305 % $\boldsymbol{\uparrow}$ 1209 % $\boldsymbol{\uparrow}$ 507 % $\boldsymbol{\uparrow}$ 1746 % $\boldsymbol{\uparrow}$ 40 % $\boldsymbol{\downarrow}$5 % $\boldsymbol{\uparrow}$ 81 %
41 × 41 $\boldsymbol{\uparrow}$ 966 % $\boldsymbol{\uparrow}$ 356 % $\boldsymbol{\uparrow}$ 1463 % $\boldsymbol{\uparrow}$ 1418 % $\boldsymbol{\uparrow}$ 541 % $\boldsymbol{\uparrow}$ 2068 % $\boldsymbol{\uparrow}$ 46 % $\boldsymbol{\downarrow}$26 % $\boldsymbol{\uparrow}$ 90 %
49 × 49 $\boldsymbol{\uparrow}$ 906 % $\boldsymbol{\uparrow}$ 345 % $\boldsymbol{\uparrow}$ 1299 % $\boldsymbol{\uparrow}$ 1719 % $\boldsymbol{\uparrow}$ 673 % $\boldsymbol{\uparrow}$ 2524 % $\boldsymbol{\uparrow}$ 86 % $\boldsymbol{\downarrow}$9 % $\boldsymbol{\uparrow}$ 144 %
63 × 63 $\boldsymbol{\uparrow}$ 993 % $\boldsymbol{\uparrow}$ 384 % $\boldsymbol{\uparrow}$ 1406 % $\boldsymbol{\uparrow}$ 2227 % $\boldsymbol{\uparrow}$ 898 % $\boldsymbol{\uparrow}$ 3137 % $\boldsymbol{\uparrow}$ 119 % $\boldsymbol{\uparrow}$ 8 % $\boldsymbol{\uparrow}$ 192 %
All $\boldsymbol{\uparrow}$ 272 % $\boldsymbol{\downarrow}$58 % $\boldsymbol{\uparrow}$ 1463 % $\boldsymbol{\uparrow}$ 381 % $\boldsymbol{\downarrow}$67 % $\boldsymbol{\uparrow}$ 3137 % $\boldsymbol{\uparrow}$ 37 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 436 %
3-dimensional neighborhood attention
3 × 3 × 3 - - - $\boldsymbol{\uparrow}$ 342 % $\boldsymbol{\uparrow}$ 110 % $\boldsymbol{\uparrow}$ 1214 % - - -
5 × 5 × 5 - - - $\boldsymbol{\uparrow}$ 244 % $\boldsymbol{\downarrow}$36 % $\boldsymbol{\uparrow}$ 1290 % - - -
7 × 7 × 7 - - - $\boldsymbol{\uparrow}$ 525 % $\boldsymbol{\uparrow}$ 20 % $\boldsymbol{\uparrow}$ 2418 % - - -
9 × 9 × 9 - - - $\boldsymbol{\uparrow}$ 466 % $\boldsymbol{\uparrow}$ 121 % $\boldsymbol{\uparrow}$ 1049 % - - -
11 × 11 × 11 - - - $\boldsymbol{\uparrow}$ 653 % $\boldsymbol{\uparrow}$ 128 % $\boldsymbol{\uparrow}$ 1511 % - - -
13 × 13 × 13 - - - $\boldsymbol{\uparrow}$ 913 % $\boldsymbol{\uparrow}$ 208 % $\boldsymbol{\uparrow}$ 2064 % - - -
15 × 15 × 15 - - - $\boldsymbol{\uparrow}$ 1243 % $\boldsymbol{\uparrow}$ 224 % $\boldsymbol{\uparrow}$ 2781 % - - -
17 × 17 × 17 - - - $\boldsymbol{\uparrow}$ 660 % $\boldsymbol{\uparrow}$ 386 % $\boldsymbol{\uparrow}$ 1037 % - - -
All - - - $\boldsymbol{\uparrow}$ 460 % $\boldsymbol{\downarrow}$36 % $\boldsymbol{\uparrow}$ 2781 % - - -
Table 4: Breakdown of improvement over naive kernels (FP32/TF32). Both GEMM-based and fused NA improve the baseline naive kernels on average. However, as illustrated by the minimum improvement, there are some cases in which naive kernels may be preferable to GEMM-based, which fails to effectively hide data prefetching latency. Fused NA on the other hand manages to improve upon most cases, with the exception of many cases that target the tiled kernels in NATTEN (2-dimensional, with window size up to 13 × 13.) Very few cases ($\approx$ 0.2 %) in 3-dimensional problems where fused NA can’t always outperform naive.
Model # of FLOPs Throughput Top-1
Params Naive GEMM Fused Accuracy
(M) (G) (imgs/sec) (%)
NAT-M 20 2.7 2975 2660 ( $\boldsymbol{\downarrow}$11 % ) 3575 ( $\boldsymbol{\uparrow}$ 20 % ) 81.8
DiNAT-M 20 2.7 2672 2548 ( $\boldsymbol{\downarrow}$5 % ) 3781 ( $\boldsymbol{\uparrow}$ 42 % ) 81.8
NAT-T 28 4.3 2167 1939 ( $\boldsymbol{\downarrow}$11 % ) 2661 ( $\boldsymbol{\uparrow}$ 23 % ) 83.2
DiNAT-T 28 4.3 1910 1845 ( $\boldsymbol{\downarrow}$3 % ) 2818 ( $\boldsymbol{\uparrow}$ 48 % ) 82.7
NAT-S 51 7.8 1457 1309 ( $\boldsymbol{\downarrow}$10 % ) 1804 ( $\boldsymbol{\uparrow}$ 24 % ) 83.7
DiNAT-S 51 7.8 1360 1313 ( $\boldsymbol{\downarrow}$3 % ) 2070 ( $\boldsymbol{\uparrow}$ 52 % ) 83.8
NAT-B 90 13.7 1110 997 ( $\boldsymbol{\downarrow}$10 % ) 1384 ( $\boldsymbol{\uparrow}$ 25 % ) 84.3
DiNAT-B 90 13.7 982 950 ( $\boldsymbol{\downarrow}$3 % ) 1463 ( $\boldsymbol{\uparrow}$ 49 % ) 84.4
DiNAT-L 200 30.6 669 647 ( $\boldsymbol{\downarrow}$3 % ) 1006 ( $\boldsymbol{\uparrow}$ 50 % ) 86.6
DiNAT-L(384 × 384) 200 92.4 153 134 ( $\boldsymbol{\downarrow}$12 % ) 301 ( $\boldsymbol{\uparrow}$ 97 % ) 87.5
Table 5: Model-level throughput changes when using our proposed GEMM-based and fused kernels in ImageNet classification. Hierarchical vision transformers NAT and DiNAT can see between 25% to 106% improvement in FP16 throughput on an A100 (batch size 128) with our proposed fused kernel. Suffering from the memory alignment issue, our half precision GEMM kernels usually result in a much smaller improvement over naive kernels, particularly the tiled variants. The same measurements with FP32 precision are presented in Tab. 6.
Model # of FLOPs Throughput Top-1
Params Naive GEMM Fused Accuracy
(M) (G) (imgs/sec) (%)
NAT-M 20 2.7 2416 2481 ( $\boldsymbol{\uparrow}$ 3 % ) 2408 ( 0 % ) 81.8
DiNAT-M 20 2.7 2217 2364 ( $\boldsymbol{\uparrow}$ 7 % ) 2690 ( $\boldsymbol{\uparrow}$ 21 % ) 81.8
NAT-T 28 4.3 1739 1802 ( $\boldsymbol{\uparrow}$ 4 % ) 1776 ( $\boldsymbol{\uparrow}$ 2 % ) 83.2
DiNAT-T 28 4.3 1591 1706 ( $\boldsymbol{\uparrow}$ 7 % ) 1988 ( $\boldsymbol{\uparrow}$ 25 % ) 82.7
NAT-S 51 7.8 1160 1199 ( $\boldsymbol{\uparrow}$ 3 % ) 1185 ( $\boldsymbol{\uparrow}$ 2 % ) 83.7
DiNAT-S 51 7.8 1102 1183 ( $\boldsymbol{\uparrow}$ 7 % ) 1389 ( $\boldsymbol{\uparrow}$ 26 % ) 83.8
NAT-B 90 13.7 867 897 ( $\boldsymbol{\uparrow}$ 3 % ) 890 ( $\boldsymbol{\uparrow}$ 3 % ) 84.3
DiNAT-B 90 13.7 795 851 ( $\boldsymbol{\uparrow}$ 7 % ) 995 ( $\boldsymbol{\uparrow}$ 25 % ) 84.4
DiNAT-L 200 30.6 506 540 ( $\boldsymbol{\uparrow}$ 7 % ) 629 ( $\boldsymbol{\uparrow}$ 24 % ) 86.6
DiNAT-L(384 × 384) 200 92.4 116 115 ( $\boldsymbol{\downarrow}$1 % ) 162 ( $\boldsymbol{\uparrow}$ 40 % ) 87.5
Table 6: Model-level throughput changes when using our proposed GEMM-based and fused kernels in ImageNet classification (full precision). While fused attention kernels are not expected to have as large of an edge over BMM-style attention kernels in FP32, our fused kernels still happen to outperform naive kernels in full precision. It is also visible that our GEMM kernels can outperform naive kernels when we eliminate the memory alignment issue. That said, our FP32 GEMM kernels still impose a maximum alignment of 1 element on the attention weights tensor, which limits its ability to compete with other BMM-style attention kernels.
Dataset # of Throughput FID
Params Naive GEMM Fused
(imgs/sec)
FFHQ (256 × 256) 48.9 M 36.7 40.6 ( $\boldsymbol{\uparrow}$ 11 % ) 44.2 ( $\boldsymbol{\uparrow}$ 21 % ) 2.05
FFHQ (1024 × 1024) 49.4 M 8.2 8.5 ( $\boldsymbol{\uparrow}$ 3 % ) 10.4 ( $\boldsymbol{\uparrow}$ 26 % ) 4.17
Table 7: Model-level throughput changes when using our proposed GEMM-based and fused kernels in style-based image generation. We benchmark StyleNAT [21], a style-based generative adversarial model based on neighborhood attention under different kernels. We experimented with different batch sizes in order to achieve peak performance, and settled for 64 for the 256 × 256 variant, and 8 for the 1024 × 1024. StyleNAT does not recommend lower-precision, therefore these measurements are only done in FP32.
In addition to our operation-level benchmarks, we also evaluate the effect of our proposed methodology on existing models that use neighborhood attention as a primitive, NAT [8] and DiNAT [7]. We benchmark the throughput from all variants according to ImageNet-1K [11] specifications, and report FP16 and FP32 measurements in Tab. 5 and Tab. 6 respectively. We also benchmark a style-based generative adversarial (GAN) model based on neighborhood attention, StyleNAT [21] and report performance improvements in Tab. 7. We find that at least in problem sizes that the ImageNet classification models NAT and DiNAT typically require, which are typically smaller in spatial size and window size, and larger in batch size, our GEMM-based approach fails to improve the baseline in half precision, and only minimally improves it in full precision. Our fused kernels on the other hand never fail to improve upon the baseline, but they only provide significant improvement in half precision, and cases that use dilation frequently (DiNAT [7] variants). Improvements in the generative model, StyleGAN [21], are only observed in full precision (half precision is not recommended in this application), where again we find that both our GEMM-based and fused kernels can improve inference speed compared to existing naive kernels, with our fused kernels having a much more noticeable edge.
5 Conclusion
In this work, we formulated the neighborhood attention problem, and by extension the sliding window attention problem, in both single-rank (1-D) and higher-rank spaces (2-D and 3-D), which are inherently GEMV problems as GEMM problems. Through this finding, we implemented extensible GEMM-based and fused CUDA kernels that implement neighborhood attention, which can significantly improve upon existing kernels in the $\mathcal{N}ATTEN$ project. These kernels will not only speed up previously-proposed models based on neighborhood attention, but can also significantly enhance ongoing research efforts in this direction. In addition, our fused kernels are at the same time the most flexible in terms of parameterization (support for varying window size, dilation factor, and causal masking across different ranks) and reduce the memory footprint of unfused (BMM-style) implementations. We however note that this is an ongoing effort, which does not yet include support for back-propagation, but at the same time we see no obstacles in extending this implementation to support backward passes, as well as additional features.
Acknowledgements. We would like to thank NVIDIA and members of the CUTLASS project [18], in particular Haicheng Wu, for his valuable feedback and comments which led to the creation of GEMM-based NA. We also thank Meta xFormers team for developing FMHA, which is what our fused neighborhood attention kernels are based on. A. H. thanks Bing Xu, Hao Lu, and Michael Iovine for the invaluable learning experience while interning at HippoML. A. H. also thanks Kai Wang for his feedback on the paper. This research was supported in part by the Intelligence Advanced Research Projects Activity (IARPA) under Contract No. 2022-21102100004, and cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA.